OpenStudy (anonymous):

ʃ(x+4/x^2+5x+6) evaluate

OpenStudy (anonymous):

=ʃ(x+4x^(-2)+5x+6 dx =(x^2)/2+4(x^(-1))/(-1)+5(x^2)/2+6x+C =1/2(x^2)-4/x+(5/2)x^2+6x+C

OpenStudy (anonymous):

thank you I was confused with your format but at the end i think it was clear what you meant. = ʃ(x+4)(x^(-2) +5x+6)dx =(x^2 /2)+(4x^(-1) /(-1))+(5x^2 /2)+6x+c =(1/2)x^2 -(4/x)+(5/2)x^2 +6x+c correct?

OpenStudy (anonymous):

yeap, that's what I meant.